OpenStudy (anonymous):

In which quadrant does the terminal side of a 118º angle lie?

OpenStudy (tkhunny):

This is a memorization problem. \(0º \lt \theta \lt 90º\) Quadrant I \(90º \lt \theta \lt 180º\) Quadrant II \(180º \lt \theta \lt 270º\) Quadrant III \(270º \lt \theta \lt 360º\) Quadrant IV Just get it stuck in your head.